PRODUCTS OF BLOCK TOEPLITZ OPERATORS

CAIXING GU DECHAO ZHENG

Abstract. In this paper we characterize when the product of two block Toeplitz operators is a compact perturbation of a block Toeplitz operator on the Hardy space of the open unit disk. Necessary and sufficient conditions are given for the commutator of two block Toeplitz operators to be compact.